OpenStudy (anonymous):

Help! I have 3 questions! The correct simplification of the expression 3b/a^-2? The correct simplification of the expression (-9a^3b^2c^10)^0? Choose the correct simplification of the expression -6x^2(4x - 7x^2 - 2)?

OpenStudy (accessdenied):

1) A negative exponent can be made positive when taking its reciprocal 1/a^(-2) = a^2/1 Simplest term means to remove those negative exponents 2) Anything to the 0 power is 1. Consider that and the exponent properties (abc)^d = a^d b^d c^d. 3) I'm guessing that you multiply the outer term through, but it sounds like there's a list to choose from...

OpenStudy (anonymous):

oh here is the list -13x^4 - 2x^3 - 8x^2 42x^4 - 24x^3 + 12x^2 42x^4 + 24x^3 - 12x^2 -42x^4 + 24x^3 - 12x^2

OpenStudy (accessdenied):

Okay, then yeah: -6x^2(4x - 7x^2 - 2) You can distribute the outer factor through with distributive property, and then just multiply them together. Then you reorder the terms from greatest degree to least to match up with the one on the list.

jhonyy9 (jhonyy9):

1. 3(a^2)b 2. 1 3. -24x^3 +42x^4 +12x^2

OpenStudy (anonymous):

^I think that is in the wrong order

OpenStudy (accessdenied):

yeah, it should be reordered with the fourth degree term first to match up with the one on your list, but its essentially the same expression.